### v1

# Remote electric-field control of antiferromagnetic magnon-polaritons

Abstract
The control of hybrid light-matter states, specifically magnon-polaritons that emerge from the strong coupling between magnons and cavity photons, remains a key challenge in developing reconfigurable quantum and classical devices. Here, we showcase the ability to remotely control antiferromagnetic magnon-polaritons at room temperature using electric field by integrating a highly birefringent liquid crystal layer into a terahertz Fabry–Pérot cavity containing an antiferromagnetic crystal. Positioned several millimeters from the magnetic material, the liquid crystal allows for adjusting the cavity’s photonic environment through electric field. This adjustment, in turn, influences the coupling strength of a particular cavity mode to the magnon resonance, thereby controlling the extent of magnon dressing by cavity photons. Our approach facilitates dynamic and reversible tuning of magnon-photon hybridization without the need for direct electrical contact or alterations to the magnetic medium. These findings create the conditions for voltage-programmable terahertz magnonic devices and new possibilities for noninvasive control strategies in spin-based information processing technologies.
keywords:
[…]
Magnons are quanta of collective excitation of the electron spins in the magnetically ordered crystals. The ability to transmit information without a charge transfer, thus without heat dissipation, makes them highly attractive to opto-spintronics. Antiferromagnetic magnons have drawn immense attention recently, as they exhibit dynamics in the terahertz (THz) range, much faster than in the case of their ferromagnetic counterparts characterized by the gigahertz (GHz) dynamics. A prerequisite for implementing antiferromagnetic magnons in opto-spintronics is their control by voltage bias, i.e., without electrical current flow. Such type of manipulation of magnons at ambient conditions has been shown exclusively in the case of the multiferroic material BiFeO31, 2 so far. However, the development of a versatile method for voltage-bias control of THz magnons applicable for any type of insulating antiferromagnets at room temperature has been still a challenge.
The strong coupling of light to quasiparticles, such as magnons, excitons, and plasmons, gives rise to new hybrid quasiparticles that can be controlled through their photonic components. In particular, the strong magnon-photon interaction leads to the formation of magnon-polaritons (MPs). These MPs exhibit both light- and matter-like properties, making them highly promising for next-generation high-speed information processing technologies. 3, 4 Over the past decade, cavity magnonics studies have been focused almost entirely on ferromagnetic magnons excited in the GHz range and tuned by the magnetic field.5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15 Only a few reports on modification of the GHz ferromagnetic MPs induced by the electric current16 or voltage bias17, 18 exist. The strong coupling to light of antiferromagnetic magnons in the THz range was demonstrated recently, mostly using temperature tuning.19, 20, 21, 22, 23, 24, 25, 26, 27 The magnetic field and temperature tuning of antiferromagnetic magnon-polaritons is either intrinsically inefficient or too slow from the view of practical applications. Despite being highly desired and crucial for applications in THz opto-spintronics, voltage bias manipulation of THz MPs has not yet been reported.
Liquid crystals provide a highly versatile platform for THz photonic applications, including voltage-induced THz beam manipulation28, 29, phase modulation30 and shifting,31 as well as the tunability of THz metamaterials.32, 33, 34, 35, 36, 37
In this work, we demonstrate the remote electrical control of THz MPs using a liquid crystal cell. We achieve it by integrating, into a Fabry-Pérot cavity, the liquid crystal cell and a 510-m-thick slab nickel oxide (NiO), a room-temperature antiferromagnetic insulator. A separation between the antiferromagnetic layer and the liquid crystal is approximately 2.5 mm. By varying the voltage applied to the liquid crystal, we tune the spatial overlap of the optical modes of the Fabry-Pérot cavity and the antiferromagnet layer (see Figs. Fig. 1(a) and Fig. 1(b)). This allows us to tune the interaction strength between the Fabry-Pérot cavity mode and the magnon above room temperature without external magnetic fields. The presented method of control of magnon-polaritons without direct electrical contact with the magnetic layer can be applied to a wide range of magnetic materials. As such, it is highly attractive for implementation in the THz opto-spintronic systems for control of information transfer.
1 Results and discussion
We measure the reflection from the Fabry-Pérot cavity composed of NiO and liquid crystal, separated by 2.5 mm (see Fig. 1(a)), using a THz time-domain spectrometer (TDS). We collect the spectra as a function of the NiO temperature () and the bias voltage () applied to the liquid crystal (see the Methods section for details). Two selected time-domain traces for V and V, collected at K, are presented in Fig.1(c). A clear voltage-induced change in the signal is observed around ps, corresponding to a THz pulse reflected from the liquid crystal.
The temperature dependence of reflection spectra of the studied Fabry-Pérot cavity at the voltage bias fixed at V is shown in Fig. 2. Three slightly sloped lines represent the three subsequent optical modes of the Fabry-Pérot cavity (labeled as , , and ). As a magnetic dipole, the magnon line is weakly visible in the spectrum but can manifest through its interactions with the cavity modes. Consistently, as it redshifts with increasing , a series of resonances with the optical modes is observed that enhance the magnon absorption. The strong coupling of the magnon with the middle Fabry-Pérot M2 at around K and THz is the most pronounced, as manifested by the largest anticrossing with the emergence of the lower (MP2) and the upper (MP3) magnon-polariton modes. The splitting between the branches MP2 and the MP3 attains GHz. The interaction of the magnon with the modes M1 and M2 is much weaker. We attribute the strongest coupling observed in the case of the M2 to its largest spatial overlap with the NiO layer. Our calculations of the spatial mode distribution presented in the further text confirm such an interpretation.
To determine MP coupling strength, we fit magnon-polariton energies obtained in a frame of a four-level coupled oscillator model38, 39 to the experimental data. The interaction between the magnon and the three optical modes of the Fabry-Pérot cavity is described by the Hamiltonian:
where Magnon denotes the energy of bare (uncoupled) magnon, Mi (, , ) denote energies of three selected optical modes of the Fabry-Pérot cavity, and (, , ) represent the interaction strength between the magnon and the respective optical mode. We fitted the eigenvalues of to the minima of the -dependent reflection spectra for each applied (see Fig. 2 for the spectra at V). In that way, we obtain the temperature dependencies of the polariton frequencies MPi (, , , ). The frequencies of the uncoupled magnon and Fabry-Pérot cavity modes are determined from the spectra recorded in the conditions, where the magnon-mode coupling is negligible.
To demonstrate remote electric-field control of MP coupling, we apply a voltage bias to the liquid crystal cell. An electric field causes the liquid crystal molecules to reorient. At V, the molecules in the liquid crystal are oriented along the plane of the liquid crystal cell (the configuration known as planar alignment40), as schematically shown in Fig. 1(a) and in the Methods section. When the exceeds a threshold value (), the molecules in the liquid crystal start to reorient continuously with the increasing voltage. At a saturation voltage (), the molecules are aligned perpendicular to the liquid crystal cell plane, as shown in Fig. 1(b). The and depend, in general, on the liquid crystal cell construction; in our case, V and V. The reorientation of the molecules causes a continuous change in the refractive index of the liquid crystal, from the initial value at 0 V bias to at the saturation bias. The change in the refractive index results in modification of frequencies and electromagnetic field spatial distribution of the Fabry-Pérot modes.
The temperature dependence of reflection spectra of the studied Fabry-Pérot cavity for selected values of the varied from 0 V to 30 V are shown in Fig. 3. We focus on the interaction between the magnon and the and modes characterized by the strength and , respectively. With the increasing , we observe an increase of the splitting between polariton branches MP3 and MP4 in the vicinity of the resonance of the uncoupled and magnon transitions occurring at around THz and K. The shift of the splitting center from the resonance results from the contribution from the mode to the wavefunctions of theMP3 and MP4 wavefunctions. The increase of the splitting saturates at V. On the contrary, the increase of the from 0 V to leads to a decrease in the splitting between the branches MP2 and MP3, indicating a reduction of the interaction strength of the magnon and the M2 mode. Frequencies of the uncoupled and modes are only weakly affected by the voltage bias, with the blueshifts not exceeding 5 GHz in the considered voltage range.
In Fig. 4(a), we present reflection spectra in the waterfall format for selected values of at the temperature of 353 K. We clearly see that the interaction of the M3 with the magnon results in the anticrossing and formation of MPs MP3 and MP4 marked with blue and red arrows, respectively. We fitted two Gaussian curves to the MP3 and MP4 transitions to quantify their amplitudes and full widths at half maximum (FWHM). The fits show voltage-induced transfer of amplitude from the MP4 to the MP3 (see Fig. 4(b)). At around 15 V, the resonance between the magnon and occurs, and their interaction is maximum. As expected, at resonance, the amplitudes and the linewidths (see Fig. 4(c)) of MP4 and MP3 are equal.
We determine the coupling strengths by fitting the eigenvalues of to the reflection minima as a function of . Values of and as a function of at K obtained from the fit are presented in Fig. 4(d). Notably, we observed a voltage-induced two-fold increase of and a drop of to around 0.7 of its value at 0 V. Our transfer matrix method calculations of the spatial distribution of the modes (see Fig. 4(e)) indicate that the overlap of the mode increases two times when the voltage is changed from 0 V to 30 V. In the case of the overlap decreases to around 0.65 of the value at 0 V. This confirms that the application of the voltage bias to the liquid crystal for tuning the total spatial overlap of the magnetic field (-field) of the cavity mode with the antiferromagnetic NiO layer provides an efficient means of control of the magnon - mode interaction strength.
2 Conclusions
We have shown the remote electrical control of THz magnon-polaritons in a Fabry-Pérot cavity embedding an antiferromagnetic NiO layer and a highly birefringent nematic liquid crystal. By applying the voltage to the liquid crystal layer, we tune the liquid crystal refractive index and, in consequence, the spatial distribution of the Fabry-Pérot cavity modes. The change of cavity mode -field overlap with the NiO layer tunes the magnon-photon coupling strength. Importantly, these results have been obtained above room temperature, without external magnetic fields, and with the liquid crystal layer millimeters away from the magnetic material. Reported remote electric-field tunability is a key advancement for remote control of information transfer in the THz opto-spintronics.
3 Methods
3.1 Experimental setup
The scheme of the experimental setup is presented in Fig. 5(a). We used a NiO single crystal with dimensions of 550.3 mm3, cut in the 111 plane (). The low spin-damping rates of NiO make it an excellent material for MP studies. We tune the magnon frequency by fixing the NiO crystal on a copper mirror using thermo-conductive paste. The mirror is placed on top of a stack of Peltier elements. The NiO temperature is varied in a range of – K, with a step size of K, while the voltage range is – V, with a step size of V. Crystal temperature is monitored with a -type thermocouple sensor placed in the copper mirror. Temperature is stabilized with a software-based PID loop controlling the current supplied to the stack of Peltier elements. The Peltier elements with the mirror and NiO are placed on a kinematic mount and x,y,z stage. The liquid crystal cell is placed on another kinematic mount, allowing precise positioning of the cell to NiO crystal.
We measure THz reflection spectra from the Fabry-Pérot cavity at 0 degree incidence angle. This is achieved using a transceiver antenna integrating source and detector (). We use two parabolic mirrors to focus the incident beam on the Fabry-Pérot cavity and to collect the reflected THz pulses (Fig.5(a)). The THz optical path is filled with dry air to minimize water-vapor absorption. To obtain spectra in the frequency domain, we applied a fast Fourier transform to measured time-domain traces, which are zero-padded from 200 ps to 1000 ps. The absolute value of obtained spectra is normalized to each spectra maximum.
3.2 Liquid Crystal Cell
We present a schematic of the liquid crystal cell cross-section in Fig. 5(b). We use the liquid crystal cell functionalized with a highly birefringent nematic liquid crystal mixture 1825, which is characterized by ordinary and extraordinary refractive indices of and , respectively41, 42. We realize a planar cell arrangement, where a 100 m thick layer of liquid crystal is sandwiched between two pristine slabs of fused quartz (Silica), each 1500 m thick. The surfaces of the silica slabs facing the liquid crystal are coated with a thin film of a conducting polymer Electra 92 (AR-PC 5090.02 by ) and uniformly rubbed with a 100 nm-thick PMMA resist. The PMMA film served as an alignment layer, facilitating the alignment of the liquid crystal molecules near its surface in the direction of rubbing. We achieve planar alignment of the liquid crystal in its bulk by matching the rubbing directions at the opposite sides of the cell.
The coating parameters for each layer are as follows: first, the silica glasses are covered with the Electra 92 via spin-coating (2000 rpm for 60 s) and baked on a hotplate at 105∘C for 5 minutes. According to the product information, the resulting coated film should have a thickness of 60 nm. Next, a 100 nm layer of PMMA resist (200K, 2 by ) is deposited on top of the Electra 92 layer by spin-coating (6000 rpm for 45 s). The resist is baked on a hotplate at 150∘C for 10 minutes. Electra is a conductive coating used in e-beam lithography, and it is semi-transparent in the THz range, so this layered structure (Electra-PMMA) forms a THz semi-transparent electrode.
Electrical contacts to the liquid crystal are made by wire soldering to the conducting polymer layer through the PMMA using tin. A signal generator with a signal amplifier supplies the liquid crystal with a sinusoidal signal at a frequency of about 80 Hz.
3.3 Numerical calculations
We perform qualitative modeling of electromagnetic field distributions of the Fabry-Pérot cavity modes using the transfer matrix method. We calculate the reflection from a system composed of parallel-plane slabs of silica glass, liquid crystal, NiO, and a gap filled with air between NiO and the liquid crystal cell. We neglect the thin layers of PMMA and Electra 92 in our calculations, as well as reflections from the copper mirror. We normalize the -field for 30 V for each mode to the -field at 0 V in silica glass. We assume thicknesses and refractive indices: silica 1.5 mm , liquid crystal 0.1 mm , and NiO 0.51 mm . In this analysis, we focus on modes that remain unaffected by interactions with the magnon mode, setting = 1.
4 Acknowledgements
The work was supported by the European Union through the ERC-ADVANCED grant TERAPLASM (No. 101053716). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. We acknowledge the support of the "Center for Terahertz Research and Applications (CENTERA2)" project (FENG.02.01-IP.05-T004/23) carried out within the "International Research Agendas" program of the Foundation for Polish Science co-financed by the European Union under the European Funds for a Smart Economy Programme. This work was partially supported by Pasific2 of the Polish Academy of Sciences, sponsored by the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No.847639 and from the Ministry of Education and Science of Poland. This work was also partially supported by the Sonata bis-13 2023/50/E/ST3/00584 grant of the National Science Centre of Poland. R. K. and O. S. acknowledge the financial support from the National Science Centre under grant SONATA BIS-12 No. 2022/46/E/ST7/00454.